Un-ionized ammonia (NH3) is toxic to fish at low concentrations. If the maximum allowable concentration of NH3 in a river (CNH3) is 0.20 mg/L as N, what is the maximum concentration of ammonium ion (NH4+) allowed (CNH4+), in units of mg/L as N, when the river water pH is 9? The dissociation of ammonium ion in water has an equilibrium constant of pKa = 9.25 and is described by the reaction: NH4+ = NH3 + H+. Show units throughout the calculation
Answer #1

From Hasselbalch equation for reaction

Now in question pH = 9

Concentration of Conjugate Base (NH3) = 0.20 mg/L

pKa = 9.25

Putting these values in Hasselbalch equation

9 = 9.25 + log(0.20)/[NH4+]

log (0.20)/[NH4+] = -0.25

Taking antilog from both side

(0.20)/[NH4+] = 0.5624

[NH4+] = 0.20/0.5624 = 0.355

Thus maximum allowed concentration of ammonium ion is 0.355 mg/L